OpenStudy (anonymous):

choose weather the following is always, sometimes or never true. x^3=3x a.always b.sometime cnever

OpenStudy (jamesj):

*whether ;-)

OpenStudy (anonymous):

try solving x^3 = 3x

OpenStudy (jamesj):

that's not right. There are solutions to x^3 = 3x x^3 = 3x if and only if x^3 - 3x = 0 x(x^2 - 3) = 0 Now are there solutions to that last equation?

OpenStudy (anonymous):

Sometime. square root of 3.

OpenStudy (anonymous):

Yea, sorry. My mistake.